# Interactive Navigation in Environments with Traversable Obstacles Using Large Language and Vision-Language Models

### 31 Jan 2024 v2

@@ I INTRODUCTION
Interacting with robots for navigation has been studied via many interfaces such as gestures, teleoperation devices, and natural languages. [4, 5, 6] recognized simple human gestures by using onboard cameras to control robots’ behavior like their forward-backward direction. Besides, different teleoperation devices such as joysticks [7] and tablets [8] can be used for navigating robots to desired places. We find these two kinds of interfaces may not be available for patients in bed. A more natural interface is to use voices, i.e., talking to the robot. Early works on robot navigation using natural languages relied on natural language processing techniques to obtain semantic information such as landmarks [9]. Recently, the birth of Large Language Models (LLMs) greatly improved the capability of extracting semantic information [10, 11, 12, 13]. Although LLMs have shown much better generalization, their output cannot be understood by navigation algorithms. For example, a motion planner cannot understand “curtains” extracted by LLMs. To bridge the gap, people started to combine language and vision-language large models. [14] used Vision-and-Language Models (VLMs) to associate images with texts from LLMs and then a visual navigation model is used to generate a feasible path based on the images from VLMs. Since the combination of a pre-trained language and a vision-language model can parse complex high-level instructions for robots, in this paper, we choose to use this paradigm to deal with human’s natural language in real-time interaction. In our case, we cannot use visual navigation methods to generate feasible paths because the landmarks in visual navigation methods (e.g., [14][15]) are normally assumed to be untraversable, thus no feasible path can be found when the goal (e.g., a bed) cannot be detected (see Fig. 1).
@@ II-B Framework Overview
With an action-aware costmap, many off-the-shelf path planning methods can be employed for the part of robot navigation. In this paper, we choose to use the A star ($A^{*}$) search algorithm [28]. A feasible path including a series of waypoints $\{w_{1},w_{2},\cdots,w_{m}\}$ can be obtained, which is not attainable without using action-aware costmaps (see Fig. 2 (b)).
@@ III-B Vision-Language Model
The landmarks obtained from the LLM output are fed to the VLM in the form of a simple prompt “curtain, chair”. We utilize Grounding DINO [21] to obtain the bounding box $B$ for each landmark. The final outputs of the large model module for each landmark form the input, $\{c_{x},c_{y},w,h,P_{a|\ell}\}$, for the following action-aware costmap construction. Figure 3 shows some examples. It should be noted that in the example in the lower right corner, due to the limitation of the VLM, a certain part of the object may be output as a bounding box (see the two bounding boxes of the curtain). With the help of the method in Sec. IV-B, we can also handle this situation well.
@@ 4      $T_{t}$ = $T_{t}$;
5 end if
6$\left\{\left(\ell_{1},P_{a_{1}|\ell_{1}}\right),\left(\ell_{2},P_{a_{2}|\ell_{2}}\right),...,\left(\ell_{n},P_{a_{n}|\ell_{n}}\right)\right\}$ = $LLM(T_{t})$;
@@ V-A Framework Evaluation in a Simulated Environment
In the simulation, we assumed that the action-aware attribute of the “orange wooden wall” is traversable, and our textual instruction can be “Go through the orange wooden wall”. Given a goal behind the wall, no feasible path can be planned without using the action-aware costmap, as shown in Fig. 6 (a). However, we found part of the “orange wooden wall” can be cleared in the costmap if constructing an action-aware costmap and a straightforward path that crosses the “orange wooden wall” can be obtained, as shown in Fig. 6 (b), demonstrating the effectiveness of the proposed framework.
@@ V-B Framework Evaluation in Real-world Environments
We conducted experiments in two real-world scenarios: a laboratory scenario and a medical scenario at the Multi-Scale Medical Robotics Center (MRC) [^2] https://www.mrc-cuhk.com/. We set up a simulation-like scenario in both and ran the SLAM algorithm [32] to construct a global map and annotate the goal (e.g., a bed in the scenario of MRC) in the map. For the traversable objects, We prepared 4 curtains with different colors and textures, which are common in our daily life, and we also prepared two different types of artificial grass.
@@ V-B2 Go through grass
We prepared two types of artificial grass: setaria viridis and phragmites australis. We mixed two kinds of grass and used “Go through the grass and watch out the chair.” as the textual instruction. Similar to the curtain experiments, a feasible path could be planned and the robot went through the grass using our framework, as shown in Fig. 7 (e).
@@ VI CONCLUSION AND FUTURE WORK
In the future, we plan to use a 3D LiDAR to replace the current 2D one to restore the 3D information of the environment for better sensing. To achieve more precise and robust action-aware segmentation of sensor data, we will deploy a state-of-the-art segmentation model (e.g., Segment Anything Model [33]) to our LM module to produce high-quality object masks for improving the capability of segmenting sensor data (e.g., LiDAR point clouds) to the background and objects.

### 13 Mar 2024 v3

@@ I INTRODUCTION
Interacting with robots for navigation has been studied via many interfaces such as gestures, teleoperation devices, and natural languages. [4, 5, 6] recognized simple human gestures by using onboard cameras to control robots’ behavior like their forward-backward direction. Besides, different teleoperation devices such as joysticks [7] and tablets [8] can be used for navigating robots to desired places. We find these two kinds of interfaces may not be available for patients in bed. A more natural interface is to use voices, i.e., talking to the robot. Early works on robot navigation using natural languages relied on natural language processing techniques to obtain semantic information such as landmarks [9]. Recently, the birth of Large Language Models (LLMs) greatly improved the capability of extracting semantic information [10, 11, 12, 13]. Although LLMs have shown much better generalization, their output cannot be understood by navigation algorithms. For example, a motion planner cannot understand “curtains” extracted by LLMs. To bridge the gap, people started to combine language and vision-language large models. [14] used Vision-and-Language Models (VLMs) to associate images with texts from LLMs and then a visual navigation model is used to generate a feasible path based on the images from VLMs. Since the combination of a pre-trained language and a vision-language model can parse complex high-level instructions for robots, in this paper, we choose to use this paradigm to deal with human’s natural language in real-time interaction. In our case, we cannot use visual navigation methods to generate feasible paths because the landmarks in visual navigation methods (e.g., [14, 15, 16, 17]) are normally assumed to be untraversable, thus no feasible path can be found when the goal (e.g., a bed) cannot be detected (see Fig. 1).
@@ II-B Framework Overview
With an action-aware costmap, many off-the-shelf path planning methods can be employed for the part of robot navigation. In this paper, we choose to use the A star ($A^{*}$) search algorithm [30]. A feasible path including a series of waypoints $\{w_{1},w_{2},\cdots,w_{m}\}$ can be obtained, which is not attainable without action-aware costmaps (see Fig. 2 (b)).
@@ III-B Vision-Language Model
The landmarks obtained from the LLM output are fed to the VLM in the form of a simple prompt “curtain, chair”. We utilize Grounding DINO [23] to obtain the bounding box $B$ for each landmark. Specifically, for each (Image, Text) pair, image and text features are extracted by an image backbone and a text backbone respectively, and then fed into a feature enhancer module to fuse cross-modality features. After that, cross-modality queries are selected using a language-guided query selection module and fed into a cross-modality decoder. The output queries of the last decoder layer are used to predict object boxes $B=\{c_{x},c_{y},w,h\}$ and extract corresponding phrases (e.g., curtain).
@@
Therefore, the final outputs of the large model module for each landmark form the input, $\{c_{x},c_{y},w,h,P_{a|\ell}\}$, for the following action-aware costmap construction. Figure 3 shows some examples. It should be noted that in the example in the lower right corner, due to the limitation of the VLM, a certain part of the object may be output as a bounding box (see the two bounding boxes of the curtain). With the help of the method in Sec. IV-B, we can also handle this situation well.
@@ 2 $T_{t}$ = $T_{t}$;
3 $\left\{\left(\ell_{1},P_{a_{1}|\ell_{1}}\right),\left(\ell_{2},P_{a_{2}|\ell_{2}}\right),...,\left(\ell_{n},P_{a_{n}|\ell_{n}}\right)\right\}$ = $LLM(T_{t})$;
@@ V-A Framework Evaluation in a Simulated Environment
In the simulation, we assumed that the action-aware attribute of the “orange wooden wall” is traversable and our textual instruction can be “Go through the orange wooden wall”. Given a goal behind the wall, no feasible path can be planned without the action-aware costmap, as shown in Fig. 6 (a). However, we found part of the “orange wooden wall” can be cleared in the costmap if constructing an action-aware costmap and a straightforward path that crosses “orange wooden wall” can be obtained, as shown in Fig. 6 (b), demonstrating the effectiveness of the proposed framework.
@@ V-B Framework Evaluation in Real-world Environments
We conducted experiments in two real-world scenarios: a laboratory scenario and a medical scenario at the Multi-Scale Medical Robotics Center (MRC:https: //www.mrc-cuhk.com). We set up a simulation-like scenario in both and ran the SLAM algorithm [34] to construct a global map and annotate the goal (e.g., a bed in the scenario of MRC) in the map. For the traversable objects, We prepared 4 curtains with different colors and textures, which are common in our daily life, and we also prepared two different types of artificial grass.
@@ V-B2 Go through grass
We prepared two types of artificial grass: setaria viridis and phragmites australis. We mixed two kinds of grass and used “Go through the grass and watch out the chair.” as the textual instruction. Similar to the curtain experiments, a feasible path could be planned and the robot went through the grass using our framework ( Fig. 7 (e)).
@@ VI CONCLUSION AND FUTURE WORK
In the future, we plan to use a 3D LiDAR to replace the current 2D one to restore the 3D information of the environment for better sensing. To achieve more precise and robust action-aware segmentation of sensor data, we will deploy a state-of-the-art segmentation model (e.g., Segment Anything Model [35]) to our LM module to produce high-quality object masks for improving the capability of segmenting sensor data (e.g., LiDAR point clouds) to the background and objects. Besides, we will upgrade the legged robot to a legged manipulator (e.g., [36]), allowing us to manipulate traversable objects and thus enhancing interactive navigation.